# Price Experimentation and Interference

### v3

@@ 6 Debiasing Canonical Estimators
In this section, we take advantage of our structural characterization of bias to introduce a novel approach to debiasing $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that applies as long as the platform has equally allocated units between treatment and control, i.e., $q=1/2$.
Recall (10) (11) provide the expressions for bias in terms of derivatives of demand. We now show how we can use observables to estimate this bias. First, we show that $D_{0y}(p_{0},p_{0},q)$ can be computed (approximately) using observed data. To see this, note that $D_{0y}(p_{0},p_{0},q)$ is defined by:
@@
By Assumption 6 we have that $D_{0}(p_{0},p_{0},q)=D(p_{0})$. Since we assume that the firm was pricing at $p_{0}$ before the experiment began, they have observed $D(p_{0})$ prior to the experiment. The quantity $D_{0}(p_{0},p_{1},q)$ is observed as well: it represents the control demand during the experiment. This means that in the limit as $p_{1}$ approaches $p_{0}$, observables can be used to approximately compute the derivative $D_{0y}(p_{0},p_{0},q)$.
We now note that by symmetry, we have $D_{0}(p_{0},p_{1},1/2)=D_{1}(p_{1},p_{0},1/2)$. This holds because both terms represent demand for the half of the goods that are priced at $p_{0}$, when the other half of goods are priced at $p_{1}$. Taking the derivative of both sides of this equality with respect to $p_{1}$, we obtain $D_{0y}(p_{0},p_{1},1/2)=D_{1x}(p_{1},p_{0},1/2)$, so that when $p_{1}=p_{0}$, there holds $D_{0y}(p_{0},p_{0},1/2)=D_{1x}(p_{0},p_{0},1/2)$. We have shown the following result.
@@
Thus if the firm is able to estimate $D_{0y}(p_{0},p_{0},q)$ using observables, then the right hand sides of (16) and (17) are debiased estimators of $GTE_{D}$ and $GTE_{\pi}$ respectively.
We note several limitations of this method. First, of course, the symmetry property being exploited only applies when the market is equally allocated to treatment and control. Second, the assumption that $D(p_{0})$ is observed relies on market stationarity; if the market is nonstationary, this adjusted estimator may not capture the true treatment effect during the experiment. Finally, estimation of $D_{0y}$ uses only half of the experimental data, namely, only the observed control demand, which may be a source of sample inefficiency (i.e., increased variance).
@@ 7 A Two-Sided Market Model
In this section, we illustrate the application of our results of the general revenue management setting introduced in Section 4 to a two-sided market model developed in Johari et al., (2022). That paper develops a stochastic dynamic market model with supply constraints in which customers book listings for a finite period of time. The paper also introduces a mean-field limit that greatly simplifies the analysis. We focus on the mean-field model here and describe its main elements.
@@ 7.1 Preliminaries
We begin by recapping the mean-field model formulation in Johari et al., (2022), adapted to our setting with price experiments. There is a two-sided platform evolving in continuous time with listings on the supply-side and customers on the demand side. There is a continuum of listings of mass $\rho$ that are homogenous with respect to their observable characteristics. Having a limited supply will play a key role in the analysis. We assume all listings have the same price $p$. At each time $t$, listings can be available or booked. The state of the system at time $t$ is given by $s_{t}$, which denotes the mass of available listings available at time $t$.
@@
Note that the general model in Johari et al., (2022) allows for heterogeneity in both listings and consumers. In the model we consider in this section, we make both consumers and listings homogeneous.
@@ 7.3 Experimental Designs
In this section we analyze two canonical experimental designs in platforms: randomizing listings or randomizing customers.
7.3.1 Listing-Side Randomization
We assume that randomization occurs on the listing side, where a fraction $q$ of the listings is assigned to treatment. Hence, a customer is exposed to a mix of treatment and control listings. In this experimental design, the balance equations are given by:
@@ 7.3.1 Listing-Side Randomization
7.3.2 Customer-Side Randomization
In this case we randomize on the customer side, with a fraction $q$ of customers randomized into treatment. In this case, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. The steady-state is given by:
@@ 7.3.2 Customer-Side Randomization
In this setting we have the following demand functions:
@@ 7.4 Properties of Biases with Market Balance
In this section we study structural properties of biases with a particular focus on extreme regimes of market balance. The analysis in this subsection closely follows Johari et al., (2022).
[…]
As the market becomes more demand constrained ($\lambda\rightarrow 0$), competition between customers declines resulting in less interference in a CR experiment. In the limit, there is no competition between customers and we expect the bias in the CR experiment to be zero. On the other hand, competition between listings increases resulting in more interference in a LR experiment. In the limit, we expect the bias in the LR experiment to be non-zero.
Conversely as the market becomes supply constrained ($\lambda\rightarrow\infty$), competition between listings declines resulting in less interference in a LR experiment. In the limit, there is no competition between listings and we expect the bias in the LR experiment to be zero. Competition between customers increases resulting in more interference in a CR experiment. In the limit, we expect the bias in the CR experiment to be non-zero.
The following propositions formalize the intuition about the behavior of the bias in both extremes of market balance.
Proposition 7.1
For all $p\in[c,+\infty)$, we have that:
$\lim_{\lambda\to 0}GTE_{D}(\lambda,p)/\lambda=\frac{\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}};\ \ \lim_{\lambda\to 0}GTE_{\pi}(\lambda,p)/\lambda=\frac{\rho v(p)}{\varepsilon+\rho v(p)}+\frac{(p-c)\rho v^{\prime}(p)\varepsilon}{(\varepsilon+\rho v(p))^{2}}.$
We have that $\lim_{\lambda\to 0}Bias_{\pi}^{CR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{CR}(\lambda,p)/\lambda=0$, while
$\lim_{\lambda\to 0}Bias_{\pi}^{LR}(\lambda,p)/\lambda=(p-c)\lim_{\lambda\to 0}Bias_{D}^{LR}(\lambda,p)/\lambda=-\frac{(p-c)\rho^{2}v(p)v^{\prime}(p)}{(\varepsilon+\rho v(p))^{2}}.$
Proposition 7.2
For all $p\in[c,+\infty)$, we have that: $\lim_{\lambda\to\infty}GTE_{D}(\lambda,p)=0$ and that:
$\lim_{\lambda\to\infty}GTE_{\pi}(\lambda,p)=\rho$. We have $\lim_{\lambda\to\infty}Bias_{\pi}^{LR}(\lambda,q)=(p-c)\lim_{\lambda\to\infty}Bias_{D}^{LR}(\lambda,q)=0$, while
$\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=(p-c)\lim_{\lambda\to\infty}Bias_{\pi}^{CR}(\lambda,p)=-\frac{\rho(p-c)v^{\prime}(p)}{v(p)}.$
In the limit where $\lambda\rightarrow 0$, Proposition 7.1 suggests that the CR experiment is unbiased while the LR experiment is biased. Similarly, in the limit where $\lambda\rightarrow\infty$, Proposition 7.2 suggests that the CR experiment is biased while the LR experiment is unbiased. Proofs of both propositions are presented in Appendix 11. The proofs are similar to the analysis in Johari et al., (2022) but we also leverage our characterization of biases based on cross-partial derivatives. In the next section we will further study the extent of these biases to characterize change of sign regimes.
@@
We also provide some figures to complement our limit results. In Figure 1. As $\lambda$ increases, the market becomes more supply-constrained and bias decreases as competition between listings dissipates. Figure 2 shows the normalized bias of profits for CR, and this dependence on $\lambda$ is naturally reversed. Both figures also show interesting behavior of biases with respect to the baseline price $p$. As the price $p$ changes, the sensitivity of demand to price changes as well; in particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit.
@@ 7.5 Change of Sign
An interesting immediate observation is that the regimes in which a change in sign occurs in an LR or CR experiment using the corresponding naive estimators are independent of the treatment fraction $q$, because the corresponding biases are independent of $q$. This has important implications for practice, as the treatment fraction is not a lever that can prevent the possibility that the sign of the treatment effect estimate is incorrect.
Using the structure of the model we can provide the following insightful characterizations for the change of sign regions, as follows.
@@
For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{LR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the listing-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{LR}_{\lambda}$.
@@
For all $\lambda\in(0,+\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of the customer-side experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p^{\prime}\in[c,+\infty)$ will be wrong if and only if $p^{\prime}\in P^{CR}_{\lambda}$.
Theorems 7.3 and 7.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run. The proofs are provided in Appendix 11.9 and repeatedly exploit the market balance equations for CR and LR respectively.
The right hand side of both the intervals $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ is the optimal monopoly price $p^{*}$, i.e., where $GTE_{\pi}(\lambda,p^{*})=0$. In particular, this reveals that the sets $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not comprised of unrealistic prices: if a firm is currently pricing optimally or very slightly below optimally, the sign of their canonical estimator will be wrong, regardless of their experimental setup.
It is not true, however, that the two experimental setups are equally fallible in all settings. Note that $\widehat{e_{p}}^{LR}$ is not equal to $\widehat{e_{p}}^{CR}$, meaning that the left endpoints of $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not the same, and behave quite differently based on market balance, as we illustrate through the following numerical study.
@@ 7.5.1 Numerical Results
Figure 3: Illustration of the change of sign region (white) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in black. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in grey.
@@
Consistent with Propositions 7.1 and 7.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment. At intermediate market balance regimes, we observe that the change of sign region corresponds to a single continuous interval on prices, as the theorems above suggest. We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected. We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
8 Fee Model
In this section we consider an environment in which a firm charges a fixed sales fee, as opposed to setting a price and incurring a marginal cost for each purchase. This represents the common situation–particularly in two-sided online marketplaces–where firms do not directly control the prices of the goods on their platforms, but instead collect a fixed fee for the sales of heterogeneous goods.
8.1 Specifications
Consider $N$ products, with a fixed vector of prices $\vec{p}\in\mathbb{R}^{N}$. The demand function $D:\mathbb{R}^{N}\to\mathbb{R}^{N}$, takes in a vector of prices and outputs a vector in $\mathbb{R}^{N}$ that represents the quantity demanded for each good. We assume that the platform selects a single fee for all products $f\in\mathbb{R}^{+}$; this fee represents a percentage of all money transacted that the platform will collect. We assume that this fee is passed on entirely to the customer and that the platform’s marginal cost is zero. Therefore, we can write the platform’s profit $\Pi(f,\vec{p})$ as follows: [^10] Prices may be a function of the fee, since firms may pass-through fee changes. However, since experiments tend to be short-lived we ignore these longer term calibration effects.
$\Pi(f,\vec{p})=f\vec{p}\cdot D((1+f)\vec{p}).$ (26)
Let $D^{\prime}(\cdot)$ represent the matrix of demand derivatives in price. We can then write the platform’s profit derivative in the fee as:
$\Pi_{f}(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p})+f\vec{p}\cdot\vec{p}\cdot D^{\prime}((1+f)\vec{p}).$ (27)
Define $R(f,\vec{p})$ as the total money transacted in the platform when the fee is $f$ and the prices are $\vec{p}$. We can then write $R(f,\vec{p})$ as
$R(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p}).$
Let $R_{f}(f,\vec{p})$ represent the derivative of this quantity with respect to the fee. We can then rewrite (27) as:
$\Pi_{f}(f,\vec{p})=R(f,\vec{p})+fR_{f}(f,\vec{p}).$ (28)
And rewrite (26) as:
$\Pi(f,\vec{p})=fR(f,\vec{p}).$
{assumption}
We assume that $\Pi_{f}$ exists and is well defined for all $f\in\mathbb{R}^{+}$ and all $\vec{p}\in\mathbb{R}^{N}$, or in other words, that $\Pi$ is differentiable over its domain.
When running an A/B experiment, the platform will compare a control fee $f_{0}$ to a treatment fee $f_{1}$. To keep the presentation succinct, in this section we focus only on estimation of profits. This would cause a true global treatment effect (GTE) of $\Pi(f_{1},\vec{p})-\Pi(f_{0},\vec{p})$. Thus we define the (infinitesimal) $GTE$ at the control fee $f_{0}$ as:
$GTE_{\Pi}(f_{0},\vec{p})=\Pi_{f}(f_{0},\vec{p}).$ (29)
We assume that a fraction $q$ of all subjects are put in treatment and $1-q$ into control. Note that as in the monopoly pricing example, this randomization can occur on either the customer or product. For simplicity, in product side randomization, we assume that for each of the $n$ types of products, a fraction $q$ of each is in treatment treatment. Let $\Pi_{0}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from control subjects during the experiment, and $\Pi_{1}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from treatment subjects. Let $R_{0}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for control goods, and $R_{1}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for treatment goods. We then write $\Pi_{0}(f_{0},f_{1},\vec{p},q)=f_{0}R_{0}(f_{0},f_{1},\vec{p},q)$ and $\Pi_{1}(f_{0},f_{1},\vec{p},q)=f_{1}R_{1}(f_{0},f_{1},\vec{p},q)$. We make the following assumptions for these functions, analogous to Assumption 4.2.
{assumption}
The following holds:

### v4

@@ 6 Debiasing Canonical Estimators
We take advantage of our structural characterization of bias to introduce a novel approach to debiasing $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$. The key to this approach, and a necessary step for it to be applied, is for the experimenter to equally allocate units between treatment and control, i.e., setting $q=1/2$.
Recall that equations (10) and (11) quantify bias in terms of the derivatives of demand. We will now show how we can use observables to estimate this bias. First, we show that $D_{0y}(p_{0},p_{0},q)$ can be computed (approximately) using observed data. To see this, note that $D_{0y}(p_{0},p_{0},q)$ is defined by:
@@
By Assumption 6 we have that $D_{0}(p_{0},p_{0},q)=D(p_{0})$. To make progress, we make the assumption that since the firm was pricing at $p_{0}$ before the experiment began, the pre-experiment observation of demand is a reasonable proxy for the control demand $D(p_{0})$. The quantity $D_{0}(p_{0},p_{1},q)$ is observed as well: it represents the control demand during the experiment. This means that in the limit as $p_{1}$ approaches $p_{0}$, observables can be used to approximately compute the derivative $D_{0y}(p_{0},p_{0},q)$.
We now note that by symmetry, we have $D_{0}(p_{0},p_{1},1/2)=D_{1}(p_{1},p_{0},1/2)$. This holds because both terms represent demand for the half of the goods that are priced at $p_{0}$, when the other half of goods are priced at $p_{1}$. Taking the derivative of both sides of this equality with respect to $p_{1}$, we obtain $D_{0y}(p_{0},p_{1},1/2)=D_{1x}(p_{1},p_{0},1/2)$, so that when $p_{1}=p_{0}$, there holds $D_{0y}(p_{0},p_{0},1/2)=D_{1x}(p_{0},p_{0},1/2)$. We have thus shown the following result.
@@
Thus if the firm is able to estimate $D_{0y}(p_{0},p_{0},q)$ using observables (as described above), then the right hand sides of (18) and (19) are debiased estimators of $GTE_{D}$ and $GTE_{\pi}$ respectively.
We note three limitations of this method. First, of course, the symmetry property being exploited only applies when the market is equally allocated to treatment and control. Second, the method is not sample efficient: only half the experimental samples, those from the control group, are used. Finally, our view that the platform can accurately estimate $D(p_{0})$ depends on the market remaining reasonably stationary during the experiment; if the market is nonstationary, this adjusted estimator may not capture the true treatment effect during the experiment.
@@ 7 A Two-Sided Market Model
In this section, we illustrate the application of our results from the general revenue management setting (introduced in Section 4) to the two-sided market model developed in Johari et al., (2022). That paper develops a stochastic dynamic market model with supply constraints in which customers book listings for a finite period of time. The paper also introduces a mean-field limit that greatly simplifies the analysis. We focus on the mean-field model here and describe its main elements.
@@ 7.1 Preliminaries
We begin by recapping the mean-field model formulation in Johari et al., (2022), adapted to our setting with price experiments. There is a two-sided platform evolving in continuous time with listings on the supply side and customers on the demand side. There is a continuum of listings of mass $\rho$ that are homogenous with respect to their observable characteristics. Having a limited supply will play a key role in the analysis. We assume all listings have the same price $p$. At each time $t$, listings can be available or booked. The state of the system at time $t$ is given by $s_{t}$, which denotes the mass of available listings available at time $t$.
@@
The general model in Johari et al., (2022) allows for heterogeneity in both listings and consumers. To simplify the analysis and exposition, in the model we consider in this section, we make both consumers and listings homogeneous.
@@ 7.3 Experimental Designs
In this section we analyze two canonical experimental designs in platforms: randomization across listings (LR experimentation) and randomization across customers (CR experimentation).
7.3.1 LR Experimentation
We assume that randomization occurs on the listing side, where a fraction $q$ of the listings are assigned to treatment. Hence, a customer is exposed to a mix of treatment and control listings. In this experimental design, the balance equations are given by:
@@ 7.3.1 LR Experimentation
7.3.2 CR Experimentation
In this setting, we randomize over customers, with a fraction $q$ of customers being placed into treatment, and the remainder in control. In this case, a treatment (resp., control) customer is exposed to the entire market under the treatment (resp., control) condition. The steady-state is given by:
@@ 7.3.2 CR Experimentation
Observe that the two terms on the right hand side above are rescaled versions of the control and treatment demand functions:
@@ 7.4 Properties of Biases with Market Balance
Johari et al., (2022) studies the bias of CR and LR experiments as the market becomes heavily demand constrained or supply constrained, respectively. Following the analysis of that paper, we obtain analogous insights, with proofs similar to Johari et al. (2022) but also leveraging our characterization of biases through cross-partial derivatives. In particular, as the market becomes more demand constrained ($\lambda\rightarrow 0$), competition between customers declines, resulting in less interference in a CR experiment; in the limit, the bias in a CR experiment approaches zero. On the other hand, in the same regime the competition between listings is high, resulting in more interference for an LR experiment, so in the limit the bias of an LR experiment remains positive. By contrast, as the market becomes more supply constrained ($\lambda\to\infty$), the opposite results hold: in this regime, competition between customers is stronger and competition between listings is weaker, so in the limit the bias of a CR experiment remains positive while the bias of a LR experiment approaches zero. We derive these results in Propositions 13.1 and 13.2 in Appendix 13.
@@
Below we provide two figures to complement our limit results. In Figures 1 and 2 we present bias normalized by the arrival rate $\lambda$. We normalize by $\lambda$ instead of the GTE because the GTE for profits converges to 0 around the optimal price. As $\lambda$ increases, the market becomes more supply-constrained and bias for a LR experiment decreases as competition between listings dissipates. Figure 2 shows the normalized bias of demand and profits in a CR experiment and this dependence on $\lambda$ is naturally reversed. Both figures also show interesting behavior of biases with respect to the baseline price $p$, a novel aspect of our setting relative to Johari et al., (2022). As the price $p$ changes, the sensitivity of demand to price changes as well. In particular, in both CR and LR experiments, both $D_{0y}$ and $D_{1x}$ tend to be smaller at larger prices (i.e., demand becomes more inelastic as the price becomes large); this behavior drives the shape of the bias curve with respect to price, and subsequently profit. Note in particular that LR experiments exhibit interesting behavior when prices are large: despite the fact that we nominally expect bias to increase when the market is demand constrained, and higher prices reduce demand in the market, nevertheless the bias drops when prices are high. This is because of the fact that elasticities change with price, and illustrates the nuanced nature of interference in price experiments.
@@ 7.5 Change of Sign
Interestingly, observe that the bias terms derived in Section 7.4 do not depend on the treatment fraction $q$ for either LR or CR experiments. In other words, the treatment fraction is not a lever that can be used to prevent—or even reduce—the possibility of a change of sign in the treatment effect estimator.
Using the structure of the model we can provide the following insightful characterizations for the change of sign regions.
@@
For all $\lambda\in(0,\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{LR}_{\lambda}\subseteq[c,\infty)$ such that the sign of an LR experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p\in[c,\infty)$ will be wrong if and only if $p\in P^{LR}_{\lambda}$.
@@
For all $\lambda\in(0,\infty)$, there exists a non-empty, closed and convex set (i.e., an interval) $P^{CR}_{\lambda}\subseteq[c,\infty)$ such that the sign of a CR experiment’s canonical estimator for $GTE_{\pi}(\lambda,p)$ at price $p\in[c,\infty)$ will be wrong if and only if $p\in P^{CR}_{\lambda}$.
Theorems 7.1 and 7.2 state that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$, such that the sign of the canonical estimator will be wrong, regardless of whether an LR or CR experiment is run. The proofs are provided in Appendix 15.7 and repeatedly exploit the market balance equations for CR and LR respectively.
The right hand side of both the intervals $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ is the optimal monopoly price $p^{*}(\lambda)$, i.e., where $GTE_{\pi}(\lambda,p^{*})=0$, for each $\lambda$. In particular, this reveals that the sets $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not comprised of unrealistic prices: if a firm is currently pricing optimally or very slightly below optimally, the sign of their canonical estimator will be wrong, regardless of their experimental setup.
It is not true, however, that the two experimental setups are equally fallible in all settings. Note that Condition (b) in 14 is different for LR and CR because the elasticities and demand sensitivities differ, meaning that the left endpoints of $P^{LR}_{\lambda}$ and $P^{CR}_{\lambda}$ are not the same. As expected based on the results of the previous section, they behave quite differently as a function of market balance, as we illustrate through the following numerical study.
@@ 7.5.1 Numerical Results
Figure 3: Illustration of the change of sign region (purple) in an LR-experiment (left) and CR-experiment (right), respectively, with $v(p)=e^{V-p}$, $V=5$, $c=0.5$, $\rho=1$, and $\epsilon=1$. The vertical axis is the (log of) the arrival rate, or market balance, paramter $\lambda$; the horizontal axis is the (baseline) control price $p$, at which the experiment is conducted. The region where condition (a) is not satisfied ($GTE_{\pi}<0$), so there is no change of sign, is in red. The region where condition (b) is not satisfied ($Bias_{\pi}^{LR}<GTE_{\pi}$ or $Bias_{\pi}^{CR}<GTE_{\pi}$, respectively), so there is no change of sign, is in blue.
@@
Consistent with Propositions 13.1 and 13.2, in Figure 3 we observe that the change of sign region becomes very small in the demand-constrained regime for a CR experiment and in the supply constrained regime for an LR experiment. At intermediate market balance regimes, we can more immediately observe that the change of sign region corresponds to a single continuous interval on prices, as Theorems 7.1 and 7.2 suggest. We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected. We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
7.5.2 Gap from Optimality
One important implication of Figure 3 is the impact that the change of sign region has on the net profits of firms who use canonical price experimentation methods. Consider the practical importance of the change of sign region’s borders. The right border of the change of sign region (where purple meets red), represents the optimal price for that particular market balance parameter $\lambda$. This is the point where $GTE_{\pi}=0$. The left border of the change of sign region (where blue meets purple) represents the final price a firm would reach, were it to repeatedly update it’s price in response to canonical A/B experimentation. This is the point where $\widehat{GTE}_{\pi}=0$, or equivalently, where $Bias_{\pi}=GTE_{\pi}$. Note that this price corresponds to the optimum suggested by repeated canonical A/B experiments. This is because, to the right of that price point, the expriment suggests decreasing the price, and to the left, to increase it.
The difference between these two prices (the true optimal price and the experimentally optimized price) can correspond to significant losses in overall profit. In Figure 4, we see the percent of true optimal profit that a firm will achieve if it uses canonical LR (left) or CR (right) experimentation, for a range of market balance scenarios. What we find is that the profit achieved from optimization via experimentation can be far below the optimal profit, with the impact worsening in highly demand-constrained markets (for LR experiments) or demand-constrained markets (for CR experiments). Even when markets are reasonably balanced (i.e., $\lambda\approx 1$), both LR and CR experimentation will lead to a profit that is more than 4% lower than optimal. Overall, these findings suggest that firms who use canonical A/B methods for price experimentation might be operating well below their potentially optimal profits, even if they are in well balanced markets. Furthermore, these losses can increase significantly as markets become more imbalanced, getting to $20-40\%$ of optimal.
Figure 4: Illustration of the gap from optimal profits that canonical A/B experimentation creates.
8 An Empirical Case Study
Thus far, we have shown that A/B pricing experiments can get the sign wrong, and that associated profit losses can be significant in a two-sided market model with logit demand. To align our analysis even more closely with practical scenarios, we empirically calibrate our model from the previous section using published data from several Airbnb markets, and investigate the conditions under which the sign-change phenomenon may occur in price experiments within this context. Although our model simplifies the complexities of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with real-world experiments. As will be shown in this section, our results suggest that a change of sign is indeed quite plausible in the empirical Airbnb context we study, demonstrating that these errors are not just a knife-edge issue—and thus potentially of high consequence for practitioners.
8.1 Model Calibration
In this section, we calibrate a specific version of our model from Section 7 to several Airbnb markets using the data from Farronato and Fradkin, (2022). Specifically, we consider the model from that section with $v(p)=e^{V-p}$. Note that this is a natural choice implied by a logit random utility model, where utility for the product is quasilinear in the price. For fixed $\varepsilon$, $V$, $\lambda$, and $p$, the resulting market balance equation is:
$\rho-s^{*}(\lambda,p)=\lambda\frac{s^{*}(\lambda,p)e^{V-p}}{\varepsilon+s^{*}(\lambda,p)e^{V-p}}.$ (28)
Recall that then in this model, $D(p)=\rho-s^{*}(\lambda,p)$, and for a given supplier cost $c$, its profit is $\pi(p)=(p-c)D(p)$.
In this exercise, we consider that $GTE$ corresponds to suppliers’ profits and that the experiment is on actual prices. It is worth noting that in practice, Airbnb does not actually set the price of listings, but instead sets fees. However, we have elected to use our price testing framework (in place of the fee-model detailed in Appendix 12) because the estimates made in Farronato and Fradkin, (2022) are better suited to examining price experiments. . Airbnb actively provides pricing guidance to hosts aimed at optimizing their pricing strategies and regularly conducts A/B tests to evaluate these recommendations (Deng and Le, 2023). This setting aligns closely with the price testing framework employed in this section. Furthermore, since fee changes are often passed on to consumers through adjustments to prices, this is a reasonable approach for our analysis.
The goal of our calibration process is to arrive at imputed values of the parameters that define the model: (1) the total mass of listings $\rho$; (2) the outside option value $\varepsilon$; (3) the cost $c$; (4) the value $V$; (5) the arrival rate $\lambda$; and (6) the reference price $p$, from which the platform runs a price experiment.
We now describe how we arrive at these six parameters. We leverage structural estimates made by Farronato and Fradkin, (2022) for Airbnb markets in major U.S. cities. In particular, Farronato and Fradkin, (2022) segment the Airbnb market into four tiers: economy, midscale, upscale, and luxury. They study each of these tiers across 10 cities. We refer to each pair of tier and city as a segment, and use superscript $s$ to denote a segment. For each segment $s$, Farronato and Fradkin, (2022) reports (1) an estimate of the average cost (per night booked), which we denote $\hat{c}^{(s)}$; and (2) an estimate of own price elasticity of demand, which we denote $\hat{e}_{p}^{(s)}$. [^11] These values come directly from tables E8 and E14 in the Appendix of Farronato and Fradkin, (2022).
For our empirical study, we treat each segment as its own instantiation of the market model in (28). For example, if we are currently considering midscale listings in Detroit, then $v(p)$ will represent the average consumer valuation for midscale Detroit listings at price $p$. On its part, $\varepsilon$ will represent the value of the “outside” option relative to this segment, accounting for all other Airbnb segments in that city and the option of not purchasing anything within the Airbnb marketplace (e.g., booking an hotel). Our calibration proceeds according to the following steps: